An aeroplane flies at a ground velocity (i.e. velocity relative to the ground) of 300 km/h N 30o W, in a wind blowing at a velocityof 50 km/h N 20o E. What is the velocity (speed and direction)of the plane relative to the ground? (Use a calculator and round the speed)



to the nearest km/h, and the corresponding angle to the nearest degree.)

1
Expert's answer
2022-06-01T10:30:32-0400
"\vec{v_p}=300\cos(90\degree+30\degree)\vec{i}+300\sin(90\degree+30\degree)\vec{j}"

"\vec{v_w}=50\cos(90\degree-20\degree)\vec{i}+50\sin(90\degree-20\degree)\vec{j}"

"\vec{v}=\vec{v_p}+\vec{v_w}=300\cos(120\degree)\vec{i}+300\sin(120\degree)\vec{j}"


"+50\cos(70\degree)\vec{i}+50\sin(70\degree)\vec{j}"

"|\vec{v}|^2=(300\cos(120\degree)+50\cos(70\degree))^2"

"+(300\sin(120\degree)+50\sin(70\degree))^2"

"\approx111783.6283"

"|\vec{v}|=\sqrt{|\vec{v}|^2}\approx334\ km/h"


"\tan \theta=\dfrac{300\sin(120\degree)+50\sin(70\degree)}{300\cos(120\degree)+50\cos(70\degree)}"

"\approx-2.30846"

"\theta=180\degree+\tan^{-1}(-2.30846)\approx113\degree"

"334\ km/h", N "23\degree"W